OpenStudy (anonymous):

Simplify (5x^-1/x^3)^-1

OpenStudy (anonymous):

equals to what ??

OpenStudy (anonymous):

\[(\frac{ 5x ^{-1} }{ x ^{3} })^{-1}\] Here. this should look better. And these are the options: a) x^4/5 b) -x^4/5 c) 5/x^4 d) x^3/5

OpenStudy (anonymous):

x^-1=1/x so (5/x^3)*1/x= 5/x^4 now (5/x^4)^-1 u turn this upside down so x^4/5 is the ans :p

OpenStudy (anonymous):

I still don't understand how it was solved but thanks for the answer :)